# Law of Vision Representation in MLLMs

### v1

@@ Abstract
We present the “Law of Vision Representation” in multimodal large language models (MLLMs). It reveals a strong correlation between the combination of cross-modal alignment, correspondence in vision representation, and MLLM performance. We quantify the two factors using the cross-modal Alignment and Correspondence score (AC score). Through extensive experiments involving thirteen different vision representation settings and evaluations across eight benchmarks, we find that the AC score is linearly correlated to model performance. By leveraging this relationship, we are able to identify and train the optimal vision representation only, which does not require finetuning the language model every time, resulting in a 99.7% reduction in computational cost. The code is available at https://github.com/bronyayang/Law_of_Vision_Representation_in_MLLMs.
@@ 1 Introduction
Current multimodal large language models (MLLMs) (Chen et al., 2024a; Liu et al., 2024e; d) have achieved remarkable advancements by integrating pretrained vision encoders with powerful language models (Touvron et al., 2023; Zheng et al., 2023). As one of the core components of a general MLLM, the vision representation is critical. Many researchers have utilized CLIP (Radford et al., 2021) as the primary image feature encoder, but its limitations are increasingly noticeable (Tong et al., 2024b; Geng et al., 2023; Yao et al., 2021). As a result, alternative vision representationand the combination of multiple vision encoders are being actively explored (Tong et al., 2024a; Lin et al., 2023).
@@
Despite this growing attention, the selection of vision representation has largely been empirical. Researchers typically test a set of vision representations on a specific MLLM and choose the one that yields the highest performance on benchmark tasks. This approach, however, is constrained by the number of representations tested and does not address the underlying factors that make certain feature representations perform better than others. As a result, the optimal vision representation for a specific MLLM is often determined by empirical performance rather than a deep understanding of the features that contribute to success. The question of what fundamentally makes a feature representation achieve the highest performance remains largely unanswered.
@@
To address this gap in understanding what makes a vision representation optimal for MLLMs, we propose the Law of Vision Representation in MLLMs. It aims to explain the key factors of vision representation that impact MLLM benchmarks performance. Our findings reveal that the cross-modal Alignment and Correspondence (AC) of the vision representation are strongly correlated with model performance. Specifically, an increase in the AC of the selected vision representation leads to improved model performance. To quantify this relationship, we define an AC score that measures cross-modal alignment and correspondence in vision representation. The AC score and model performance exhibit a linear relationship, with a coefficient of determination of 95.72%.
@@
Furthermore, the Law of Vision Representation guides the selection of an optimal vision representation for MLLMs. Originally, this process was extremely costly because even subtle changes in vision encoding methods—such as switching encoder types, altering image resolution, or testing feature combinations—require finetuning the language model (Lin et al., 2024). For example, using a top data-efficient MLLM pipeline with a 7B language model requires 3,840 NVIDIA A100 GPU hours to test the 10 encoders used in this study, amounting to a cost of approximately \$20,000111https://replicate.com/pricing. Testing additional encoders leads to a linear increase in cost. Moreover, the recent trend of feature combination, which often results in better performance, necessitates combinatorial testing of vision encoders. Testing all possible combinations of 10 encoders results in 1023 combinations, exponentially increasing the cost and energy consumption. This process consumes approximately 100,000 kilowatt-hours [^2] https://www.nvidia.com/content/dam/en-zz/Solutions/Data-Center/a100/pdf/nvidia-a100-datasheet.pdf, enough to drive an electric vehicle around the Earth 13 times.
@@
Thus, we are the first to propose a policy, AC policy, that selects the optimal vision representation using AC scores within the desired search space. Unlike traditional methods that rely on benchmarking performance, the AC policy enables the expansion of the search space—allowing for an increased number of vision representations to be considered—without incurring additional costs. We demonstrate that this approach enhances both accuracy and efficiency compared to randomly searching for the optimal representation. The policy successfully identifies the optimal configuration among the top three choices in 96.6% of cases, with only three language model finetuning across a 13-setting search space.
@@ 2.1 Vision for MLLMs
Recent studies have explored various vision representations in MLLMs (Beyer et al., 2024; Ge et al., 2024; Liu et al., 2024e; Wang et al., 2024; Sun et al., 2023; Luo et al., 2024). Interestingly, some findings indicate that relying solely on encoders outside of the CLIP family (Cherti et al., 2023; Zhai et al., 2023b; Li et al., 2022), such as DINOv2 (Oquab et al., 2023) and Stable Diffusion (Rombach et al., 2021), often leads to lower performance scores (Karamcheti et al., 2024; Tong et al., 2024a). However, combining features from these encoders with CLIP features—such as concatenating image embeddings in the token or channel dimension—significantly enhances performance beyond using CLIP alone (Tong et al., 2024a; b; Liu et al., 2024c; Kar et al., 2024). Researchers intuitively suggest that these additional encoders provide superior detail-oriented capabilities, but no studies have thoroughly analyzed the underlying causes of the performance change (Wei et al., 2023; Lu et al., 2024). This suggests that the attributes of an optimal vision representation remain not fully understood.
@@ 2.2 Cross-modal Alignment
Cross-modal alignment refers to the alignment between image and text feature spaces (Duan et al., 2022). This concept emerged with the introduction of text-image contrastive learning (Radford et al., 2021; Jia et al., 2021). Although current MLLMs utilize contrastively pretrained image encoders, the challenge of achieving effective alignment persists (Ye et al., 2024; Zhai et al., 2023a; Woo et al., 2024). Despite efforts to critique the limitations of CLIP family representations and explore alternative vision representations, many approaches continue to rely on contrastively pretrained encoders or adding contrastive loss without fully eliminating them (Zhang et al., 2024a; Lu et al., 2024; Tong et al., 2024a; b; Liu et al., 2024b). In our work, we point out that alignment in vision representation is essential for improved model performance and is crucial for data efficiency. Without pre-aligned vision representations, extensive data pretraining is required to achieve cross-modal alignment within the language model (Ge et al., 2024; Chen et al., 2024b; Li et al., 2024b).
@@ 2.3 Visual Correspondence
Visual correspondence is a fundamental component in computer vision, where accurate correspondences can lead to significant performance improvements in tasks, such as image detection (Xu et al., 2024; Nguyen & Meunier, 2019), visual creation (Tang et al., 2023; Zhang et al., 2024b), and MLLMs (Liu et al., 2024a), etc. Correspondences are typically categorized into semantic- and geometric-correspondences. Semantic correspondences (Zhang et al., 2024b; Min et al., 2019) involve matching points that represent the same semantic concept not necessarily representing the same instance. Geometric correspondences (Sarlin et al., 2020; Lindenberger et al., 2023), on the other hand, require matching the exact same point across images, which is often crucial for low-level vision tasks, such as pose estimation (Sarlin et al., 2020; Lindenberger et al., 2023; Zhang & Vela, 2015),and SLAM tasks, etc.
Several studies have pointed out that the CLIP family’s vision representation ”lacks visual details” (Lu et al., 2024; Tong et al., 2024b; Ye et al., 2024). We explain this observation through the concept of correspondence. Current multi-modal large language models (MLLMs) convert images into embeddings, with each embedding representing a patch of the image. Image features with high correspondence increase the similarity within internal image patches on similar semantics, thereby enabling the retrieval of more detailed information.
@@ 3 Law of Vision Representation in MLLMs
where $f$ is a linear function on second-degree polynomial transformations of $A$ and $C$.
@@ 3.2 Empirical Justification
In this section, we empirically show that AC is strongly correlated to model performance. To quantify the correlation between AC and model performance, we first propose methods to measure cross-modal alignment and correspondence within the vision representation:
  • To quantify cross-modal alignment, we aim to compare the image and text embeddings of the same concept. However, finding the same concept is difficult since it requires alignment. To address this, we use the CLIP vision embedding as a reference. We calculate the maximum cosine similarity $\textsc{S}_{\textsc{c}}$ between vector pairs from the CLIP embedding $\hat{E}$ and target vision representation embedding $E$:
    A $\textsc{A Score}=\frac{1}{n}\sum^{n}_{i=0}\ \max\limits_{u,v}\ \textsc{S}_{\textsc{c}}(\hat{E}_{i}^{(u)},E_{i}^{(v)})$ (2)
    where $n$ is the total number of image samples, and $E_{i}^{(v)}=MLP(F_{i})^{(v)}$ is the $v$-th embedding vector, resulting from the vision feature $F$ of the $i$-th image.
  • To compute the correspondence score, we extract features from paired images, resulting in $F^{i}_{s}$ and $F^{i}_{t}$ from a specific source and target pair. Given ground truth key points $\{p^{s}_{1},\dots,p^{s}_{m}\}$, we obtain the predicted key points $\{p^{t}_{1},\dots,p^{t}_{m}\}$ using the features. The correspondence score is the Percentage of Correct Keypoints (PCK) calculated using the following equation:
    C $\textsc{C Score}=\frac{1}{m}\sum^{m}_{i=0}\mathbbm{1}_{\left\|p^{t}_{j}-p^{s}_{j}\right\|_{2}<T}$ (3)
    where $T$ is a threshold defined as proportional to the bounding box size of the object instance in the image.
Finally, AC score is a second-degree polynomial transformation of the $A$ and $C$ score:
AC $\textsc{AC Score}=\sum^{2}_{i=0}\sum^{2-i}_{j=0}\beta_{ij}A^{i}C^{j}$ (4)
@@ Results.
Figure 1: $R^{2}$ values for linear regression models fitted on various scores.
We fit a simple linear regression model using 13 vision representations across 4 vision-based MLLM benchmarks. As shown in Figure 1, the average coefficient of determination ($R^{2}$) obtained is 95.72% when using the AC score of the vision representations. For comparison, we also fit models using 13 random scores, the A score alone, and the C score alone, all with polynomial transformations. The random scores and single-factor models show significantly lower correlations with performance. This result highlights the strong correlation between the AC score and MLLM performance, validating the Law of Vision Representation. Refer to Section 5.4 for details.
@@ 4 AC Policy
A general MLLM architecture consists of a frozen vision encoder followed by a trainable connector (alignment module) and the language model, similar to the setup in LLaVA (Liu et al., 2024e). To determine the optimal out of $k$ vision representations, we originally needs finetune LLM $k$ times, making the scaling of $k$ difficult. Therefore, given $N$ vision encoders, we propose AC policy, as illustrated in Figure 2, to efficiently estimate the optimal vision representation from a search space consisting of $k$ vision representation out of all $12^{N}-$ 1 possible feature combinations. We finetune only $k^{\prime}$ LLMs to obtain downstream performance, allowing $k$ to scale without significant cost, where $k^{\prime}\ll k$.
@@
Let $\mathbf{X}\in\mathbb{R}^{k\times 6}$ be the matrix containing AC scores of vision representation in the search space. We subsample $k^{\prime}$ data points from $\mathbf{X}$, denoted as $\mathbf{X}_{s}\in\mathbb{R}^{k^{\prime}\times 6}$, to serve as the input to the linear regression model:
$\mathbf{y}=\mathbf{X}_{s}\mathbf{w}+\mathbf{\epsilon}$ (5)
Here, $\mathbf{w}\in\mathbb{R}^{6}$ is the vector of model parameters, $\mathbf{\epsilon}\in\mathbb{R}^{k^{\prime}}$ is the vector of error terms, and $\mathbf{y}\in\mathbb{R}^{k^{\prime}}$ represents the downstream performance on a desired benchmark.
@@
Benchmark Number of Fintuning Recall@3
Random 12 92.04%
MMBench 4 90.1%
MME 3 90.6%
OKVQA 3 90.0%
SEED-Bench 3 86.0%
MMMU 7 96.5%
TextVQA 5 83.1%
VizWiz 3 84.6%
ScienceQA 3 96.6%
Average 3.88 89.69%
Table 1: Number of LLM finetunings required to achieve approximately 90% Recall@3 in predicting the optimal vision representation.

### v2

@@ Abstract
We introduce the “Law of Vision Representation” in multimodal large language models (MLLMs), revealing a strong correlation among cross-modal alignment, vision representation correspondence, and overall model performance. We quantify the these factors using the cross-modal Alignment and Correspondence scores. Extensive experiments across fifteen distinct vision representation settings and evaluations on eight benchmarks show that A and C scores correlate with performance following a quadratic relationship. By leveraging this relationship, we can identify and train the optimal vision representation for an MLLM, achieving a 99.7% reduction in computational cost without the need for repeated finetuning of the language model. The code is available at https://github.com/bronyayang/Law_of_Vision_Representation_in_MLLMs.
@@ 1 Introduction
Current multimodal large language models (MLLMs) (Chen et al., 2024a; Liu et al., 2024e; d) have achieved remarkable advancements by integrating pretrained vision encoders with powerful language models (Touvron et al., 2023; Zheng et al., 2023). Among the core components of a general MLLM, vision representation plays a critical role. Many researchers have utilized CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023b; Tschannen et al., 2025) as the primary image feature encoder, but their limitations are becoming increasingly noticed (Tong et al., 2024b; Geng et al., 2023; Yao et al., 2021). As a result, alternative vision representations (Tang et al., 2025) and the combination of multiple vision encoders are being actively explored (Tong et al., 2024a; Lin et al., 2023).
@@
Despite this growing attention, selection of vision representation has largely been empirical. Researchers typically test a set of vision representations on a specific MLLM and choose the one that yields the highest performance on benchmark tasks. This approach, however, is constrained by the number of representations tested and does not address the underlying factors that drive performance differences. As a result, the optimal vision representation for a specific MLLM is often determined by empirical performance rather than a deep understanding of the factors that contribute to success. The question of what fundamentally makes a feature representation achieve the highest performance remains largely unanswered.
@@
To address this gap in understanding what makes a vision representation optimal for MLLMs, we propose the Law of Vision Representation in MLLMs. It aims to explain the key factors of vision representation that impact MLLM benchmarks performance. Our findings reveal that cross-modal Alignment (A) and Correspondence (C) of the vision representation are strongly correlated with model performance.; specifically, higher A and C lead to improved performance. To quantify this relationship, we define A and C scores that measure cross-modal alignment and correspondence in vision representation. The A and C score as well as model performance exhibit a quadratic relationship, with a coefficient of determination of 94.06%.
@@
Furthermore, the Law of Vision Representation guides the selection of an optimal vision representation for MLLMs. Originally, this process was extremely costly because even subtle changes in vision encoding—such as switching encoder types, altering image resolution, or testing feature combinations—require finetuning the language model (Lin et al., 2024). For example, using a top data-efficient MLLM pipeline with a 7B language model requires 3,840 NVIDIA A100 GPU hours to test the 10 encoders, amounting to a cost of approximately \$20,000111https://replicate.com/pricing. Testing additional encoders leads to a linear increase in cost. Moreover, the recent trend of feature combination, which often results in better performance, necessitates combinatorial testing of vision encoders. Testing all possible combinations of 10 encoders results in 1023 combinations, exponentially increasing the cost and energy consumption. This process consumes approximately 100,000 kilowatt-hours [^2] https://www.nvidia.com/content/dam/en-zz/Solutions/Data-Center/a100/pdf/nvidia-a100-datasheet.pdf, enough to drive an electric vehicle around the Earth 13 times.
@@
Thus, we are the first to propose a policy, AC policy, that selects the optimal vision representation using AC scores within the desired search space. Unlike traditional methods that rely on benchmarking performance, the AC policy enables the expansion of the search space—allowing for an increased number of vision representations to be considered—without incurring additional costs. We demonstrate that this approach enhances both accuracy and efficiency compared to randomly searching for the optimal representation. The policy successfully identifies the optimal configuration among the top three choices in 96.6% of cases, with only three language model finetuning across a 15-setting search space.
Figure 1: Visualization of the Law of Vision Representation in MLLMs.
@@ 2.1 Vision for MLLMs
Recent studies have explored various vision representations in MLLMs (Beyer et al., 2024; Ge et al., 2024; Liu et al., 2024e; Wang et al., 2024b; Sun et al., 2023; Luo et al., 2024). Interestingly, some findings indicate that relying solely on encoders outside of the CLIP family (Cherti et al., 2023; Zhai et al., 2023b; Li et al., 2022), such as DINOv2 (Oquab et al., 2023) and Stable Diffusion (Rombach et al., 2021), often leads to lower performance (Karamcheti et al., 2024; Tong et al., 2024a). However, combining features from these encoders with CLIP features, such as concatenating image embeddings in the channel dimension, significantly enhances performance beyond using CLIP alone (Tong et al., 2024a; b; Liu et al., 2024c; Kar et al., 2024). Researchers intuitively suggest that these additional encoders provide superior detail-oriented capabilities, but no studies have thoroughly analyzed the underlying causes of the performance change (Wei et al., 2023; Lu et al., 2024a). This suggests that the attributes of an optimal vision representation remain not fully understood.
@@ 2.2 Cross-modal Alignment
Cross-modal alignment refers to the alignment between image and text feature spaces (Duan et al., 2022). This concept emerged with the introduction of text-image contrastive learning (Radford et al., 2021; Jia et al., 2021). Although current MLLMs utilize contrastively pretrained image encoders, the challenge of achieving effective alignment persists (Ye et al., 2024; Zhai et al., 2023a; Woo et al., 2024). Despite efforts to critique the limitations of CLIP family representations and explore alternative vision representations, many approaches continue to rely on contrastively pretrained encoders or adding contrastive loss without fully eliminating them (Zhang et al., 2024b; Lu et al., 2024a; Tong et al., 2024a; b; Liu et al., 2024b). In our work, we point out that alignment in vision representation is essential for improved model performance and is crucial for data efficiency. Without pre-aligned vision representations, extensive data pretraining is required to achieve cross-modal alignment within the language model (Ge et al., 2024; Chen et al., 2024b; Li et al., 2024c).
@@ 2.3 Visual Correspondence
Visual correspondence is a fundamental component in computer vision, where accurate correspondences can lead to significant performance improvements in tasks, such as image detection (Xu et al., 2024; Nguyen & Meunier, 2019), visual creation (Tang et al., 2023; Zhang et al., 2024c), and MLLMs (Liu et al., 2024a), etc. Correspondences are typically categorized into semantic- and geometric-correspondences. Semantic correspondences (Zhang et al., 2024c; Min et al., 2019) involve matching points that represent the same semantic concept not necessarily representing the same instance. Geometric correspondences (Sarlin et al., 2020; Lindenberger et al., 2023), on the other hand, require matching the exact same point across images, which is often crucial for low-level vision tasks, such as pose estimation (Sarlin et al., 2020; Lindenberger et al., 2023; Zhang & Vela, 2015), and SLAM tasks, etc.
Several studies have pointed out that the CLIP family’s vision representation ”lacks visual details” (Lu et al., 2024a; Tong et al., 2024b; Ye et al., 2024). We explain this observation through the concept of correspondence. Current MLLMs convert images into embeddings, with each embedding representing a patch of the image. Image features with high correspondence increase the similarity within internal image patches on similar semantics, thereby enabling the retrieval of more detailed information.
@@ 3 Law of Vision Representation in MLLMs
where $f$ is a quadratic function of $A$ and $C$.
3.1 Assumptions
Following NVLM (Dai et al., 2024), we categorize MLLMs into the following types: (1) Decoder-only MLLMs (Tong et al., 2024a; Liu et al., 2024e; Li et al., 2024a; Liu et al., 2024f; Dai et al., 2024; Lu et al., 2024b; Zhang et al., 2024a; Wang et al., 2024a): These MLLMs consist of vision encoder(s) and an alignment module, such as a multilayer perceptron (MLP), which maps the vision representation into vision tokens. These tokens are designed to have a similar distribution as language tokens and are directly input into a language model in the same manner as language tokens. (2) Cross-attention-based MLLMs (Dai et al., 2024; Bai et al., 2023; Alayrac et al., 2022; Laurençon et al., 2024; Chen et al., 2024c): These MLLMs include vision encoder(s) and an additional module, often serving as a downsampling component, such as a perceiver resampler. The vision tokens generated are integrated into the language model through cross-attention mechanisms.
  • The Law of Vision Representation specifically focuses on decoder-only MLLM architecture due to their widespread adoption and their simplicity, which facilitates controlling variables in training recipes and enables clear mathematical modeling.
  • We further assume vision representation is the only independent variable, while the alignment module and LLM architecture remain fixed. In the case of a unfrozen vision encoder, we cannot guarantee that the vision encoder does not take the function of the alignment module. This causes the architecture and role of the alignment module to change alongside the encoder, making the experiment uncontrolled and the models no longer comparable.
@@ 3.3 Empirical Justification
In this section, we empirically show that $A$ and $C$ scores are strongly correlated to model performance. To quantify the correlation between $A$ and $C$ as well as model performance, we first propose methods to measure cross-modal alignment and correspondence within the vision representation:
  • To quantify cross-modal alignment, we define a metric A Score, that measures how well the vision representation is mapped into the language model’s space. With both the vision encoder and the LLM frozen, if visual features aligns with the LLM’s language space effectively, the LLM’s prediction error will be minimized. In other words, a well-aligned vision embedding leads to a higher likelihood for the correct caption tokens.
    Formally, for an input image $I$ and its associated caption $y=(y_{1},y_{2},\dots,y_{T})$, where $T$ is the sequence length, let $f(I)$ denote the projected visual representation (i.e., the output of the vision encoder + projector). The conditional probability of generating token $y_{t}$ is given by:
    $P\big{(}y_{t}\mid f(I),y_{<t}\big{)},$
    , with $y_{<t}$ representing the tokens preceding $y_{t}$.
    The alignment score is then defined as the average log likelihood over all tokens:
    A $\textsc{A Score}(I,y)=\frac{1}{T}\sum_{t=1}^{T}\log P\big{(}y_{t}\mid f(I),y_{<t}\big{)}$
    A higher A Score indicates that the visual features are more effectively aligned with the language model’s representation, as reflected by the increased log-likelihood of the correct caption.
  • To quantify correspondence, we measure how accurately key points in one image can be matched to their semantically corresponding locations in another image. Given a pair of image with annotated, semantically matching key points, we first extract features from each image pair. Let $F^{s}$ and $F^{t}$ denote the feature maps of the source and target images, respectively.
    Using the feature vectors at the labeled key point positions in $F^{s}$, we predict the corresponding key points in $F^{t}$ by selecting the location with the maximum similarity, yielding a set of predicted key points $\{p^{pred}_{1},\dots,p^{pred}_{m}\}$ for $m$ key points. The ground-truth key points for the image pair are denoted by $\{p^{GT}_{1},\dots,p^{GT}_{m}\}$.
    The correspondence score is then defined as the Percentage of Correct Keypoints (PCK), computed as follows:
    C $\textsc{C Score}=\frac{1}{m}\sum^{m}_{i=0}\mathbbm{1}_{\left\|p^{pred}_{i}-p^{GT}_{i}\right\|_{2}<T}$ (2)
    where $T$ is a threshold proportional to the bounding box size of the object in the image, and $\mathbbm{1}(\cdot)$ is the indicator function that returns 1 when the condition is satisfied and 0 otherwise.
    A higher C Score indicates more accurate key point correspondence, reflecting better vision feature matching.
To capture the overall performance, we integrate the A Score and the C Score into a single metric called the AC Score. We do this by fitting a second-degree polynomial to benchmark performance, which allows us to model potential nonlinear interactions between $A$ and $C$. Formally, the AC Score is defined as:
AC $\textsc{AC Score}=\sum^{2}_{\alpha=0}\sum^{2-\alpha}_{\beta=0}w_{\alpha\beta}A^{\alpha}C^{\beta}$ (3)
where $w_{\alpha\beta}$ are trainable parameters that are optimized to best fit the benchmark performance.
This formulation allows the AC Score to capture both the individual contributions of alignment and correspondence, as well as their interaction effects.
@@ Results.
Figure 2: $R^{2}$ values for regression models fitted on various scores.
We fit a simple regression model using 15 vision representations across 4 vision-based MLLM benchmarks. As shown in Figure 2, the average coefficient of determination ($R^{2}$) obtained is 94.06% when using the AC score of the vision representations. For comparison, we also fit models using 15 random scores, the A score alone, and the C score alone, all with quadratic functions. The random scores and single-factor models show lower correlations with performance. This result highlights the strong correlation between the AC score and MLLM performance, validating the Law of Vision Representation. Refer to Section 5.4 for details.
@@ 4 AC Policy
The MLLM architecture assumed in this framework consists of a frozen vision encoder, followed by a trainable connector (alignment module) and the pretrained language model. To determine the optimal out of $k$ vision representations for the MLLM, we originally needs finetune LLM $k$ times, making the scaling of $k$ difficult. Therefore, we propose AC policy, as illustrated in Figure 3, to efficiently estimate the optimal vision representation from a search space consisting of $k$ vision representations. We finetune only $k^{\prime}$ LLMs to obtain downstream performance, allowing $k$ to scale without significant cost, where $k^{\prime}\ll k$. The value of $k^{\prime}$ should be determined based on the computational budget allocated for vision representation selection.
@@
Let $\mathbf{X}\in\mathbb{R}^{k\times 6}$ be the matrix containing AC scores of vision representation in the search space. We subsample $k^{\prime}$ data points from $\mathbf{X}$, denoted as $\mathbf{X}_{s}\in\mathbb{R}^{k^{\prime}\times 6}$, to serve as the input to the regression model:
$\mathbf{y}=\mathbf{X}_{s}\mathbf{w}+\epsilon$ (4)
Here, $\mathbf{w}\in\mathbb{R}^{6}$ is the vector of model parameters, $\epsilon\in\mathbb{R}^{k^{\prime}}$ is the vector of error terms, and $\mathbf{y}\in\mathbb{R}^{k^{\prime}}$ represents the downstream performance on a desired benchmark.